OpenStudy (anonymous):

Find the range of (x^2-9)/(x-3)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[(x-3) (x+3)/(x-3) \]

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

y = (x+3)

OpenStudy (anonymous):

cancel and get...

OpenStudy (anonymous):

then wat to do

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

make sure to recall that in your original function you cannot put \(x=3\)

OpenStudy (anonymous):

the range of fun is R - {3}

OpenStudy (anonymous):

so take \(f(x)=x+3\) replace \(x\) by 3 and get \(f(3)=6\) but \(x\) cannot be 3, so the range does not include 6

OpenStudy (anonymous):

so f(3) = 3+3 = 6 Range is R-{6}

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thxx a lot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now i got some idea to find range